The ridioisotope iodine - 131 (half-life = 8 days) is received from a medical supplier in a solution such that 0.5 cc of the solution has enough activity for injection for a certain organ scan at the time of supply. If it is impractical to use more than 8 cc of the solution per scan, what is the useful life of the solution?
Answer #1

T1/2 = half life = 8 days

\lambda = decay constant

decay constant is given as

\lambda = 0.693/T1/2

\lambda = 0.693/8 = 0.087

N = 0.5

No = 8

t = useful life

using the equation

N = No e-\lambdat

0.5 = 8 e-0.087t

t = 31.86 days